zarkam21 (zarkam21):

What is the sum of the interior angles of a 16-sided polygon?

OpenStudy (roxy.girl):

So the sum of all of the measures of the interior angles of a 16-sided polygon is 2520 degrees. S = 180(n-2) S = 180(16-2) S = 180(14) S = 2520